Mike & Bill row their boat (constant speed) 55mi downstream for 2.5hrs, helped by the current. Rowing at the same rate, the trip back against the current takes 5.5hrs. Find the rate of the current.
Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Let R=rate of boat in still water
and let r=rate of current
(DOWNSTREAM WE ADD THE RATE OF THE CURRENT; UPSTREAM, WE SUBTRACT)
Downstream:
R+r=55/(2.5)=22-------------------eq1
Upstream:
R-r=55/(5.5)=10------------------eq2
add eq1 and eq2:
2R=32
R=16 mph---speed of boat ---substitute into eq1 and we get:
r=22-16=6 mph---speed of current
CK
2.5*22=55
5.5*10=55
Hope this helps--ptaylor